An athlete at the gym holds a 3.0 kg steel ball in his hand. His arm is 70 cm long and has a mass of 4.0 kg . What is the magnitude of the torque about his shoulder if he holds his arm straight, but 30 ∘ below horizontal?
Answer #1

Solution)

Given,

Mass of steel ball, m= 3kg

Mass of arm, M= 4kg

Length of arm, l=0.70 m

So, d= l/2=0.35 m

Theta= 30 degrees

Now, we know

Torque= Mgd cos theta + mgl cos theta

Torque= gcos theta(Md + ml)

Torque= 9.8* cos 30(4*0.35 + 3*0.7)

Torque= 29.7 Nm

===================